I need help finding the distance between two linear pairs. (-1,4) and (2,-3) This is what ive tried to do but it's not right.
(4-3)^2+(-2-5)^2
1^2+-3^2
sqrt (-8)

Answer by stanbon(75887)

I need help finding the distance between two linear pairs. (-1,4) and (2,-3) This is what ive tried to do but it's not right.
-------------------
d^2 = (4-3)^2+(-2-5)^2
d^2 = (1)^2 + (-7)^2
d^2 = 1 + 49
d^2 = 50
d = sqrt(50)
d = sqrt(25*2)
d = 5sqrt(2)
=============
Cheers,
Stan H.
=============